OpenStudy (anonymous):

if cot(x)=-13/12 where x is in Q 4 what is sin(2x)?

OpenStudy (anonymous):

Are you allowed to use a calculator?

OpenStudy (anonymous):

Write it in terms of tangent:\[\Large \tan(x)=\frac{1}{\cot(x)}=\frac{1}{-13/12}=-\frac{12}{13}\]Take the arctangent (inverse tangent) of both sides.\[\Large x=\arctan(-\frac{12}{13})\approx-42.7\]Now you can find sin(2x).\[\Large \sin(2x)=\sin(2*(-42.7))=\sin(-85.4)=-0.997\approx-1\]So it's about negative 1.

OpenStudy (anonymous):

thank you